OpenStudy (anonymous):

**************RECEIVE FAN + MEDAL FOR THIS QUESTION****************************** line WX contains (-1, 2) and (4 ,12) line YZ contains points (5, 8) and (-2, -6) . Lines WX and YZ are.

OpenStudy (anonymous):

Perpendicular, because the slopes are the same Parallel, because the product of the slopes is -1 Perpendicular, because the product of the slopes is -1 Parallel, because the slopes are the same

OpenStudy (anonymous):

parallel because the slopes are the same the slope for WX: \[\frac{ 12-2 }{ 4-(-1) }=2\] the slope for YZ: \[\frac{ -6-8 }{ -2-5 }=2\] So they are both equal thus they are parallel

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

:)